# Analyzing Cognitive Differences Among Large Language Models through the Lens of Social Worldview

### v1

@@ H.1 Statistical Methodology
The repeated-measures ANOVA model is formally expressed as:
where $\mu$ denotes the overall grand mean; $\alpha_{j}$ captures the main effect of the cognitive dimension; $\beta_{k}$ is the main effect associated with the experimental condition; $(\alpha\beta)_{jk}$ is the dimension-condition interaction; $s_{i}\sim N(0,\sigma_{s}^{2})$ represents the random effect for LLM-specific variability; and $\varepsilon_{ijk}\sim N(0,\sigma^{2})$ denotes the residual error term.
We first tested the global null hypothesis of no difference among experimental condition means:
$\begin{split}H_{0}:\quad\beta_{\text{Awareness}}&=\beta_{\text{Feedback-None}}\\ =\beta_{\text{Feedback-Less}}&=\beta_{\text{Feedback-Most}}\end{split}$
Upon rejecting this null hypothesis at significance level $p<0.05$, we proceeded with Bonferroni-adjusted pairwise post-hoc comparisons between the Social Self-Awareness condition and each distinct feedback condition. The paired differences for each dimension $j$ and each feedback condition were computed as:
$D_{ij}^{(\text{Awareness\textendash Feedback})}=Y_{ij,\text{Awareness}}-Y_{ij,\text{Feedback}},$
and evaluated via paired $t$-statistics given by:
$t_{j}$ $=\frac{\bar{D}_{j}}{s_{D_{j}}/\sqrt{N}},\quad\text{where}$
$\bar{D}_{j}$ $=\frac{1}{N}\sum_{i=1}^{N}D_{ij}^{(\text{Awareness\textendash Feedback})},$
$s_{D_{j}}$ $=\sqrt{\frac{\sum_{i=1}^{N}(D_{ij}^{(\text{Awareness\textendash Feedback})}-\bar{D}_{j})^{2}}{N-1}}.$
We quantified practical significance using partial eta-squared ($\eta_{p}^{2}$) from the ANOVA results, calculated as:
and the standardized mean difference (Hedges’ $g$) for each pairwise post-hoc test:
Polynomial Trend Analysis of Feedback Intensity (RQ3b).
To characterize how graded intensities of social feedback incrementally influence LLM cognitive outcomes beyond the effect of social self-awareness, we conducted a polynomial trend analysis on the difference scores between each feedback condition and the baseline Self-Awareness condition. Specifically, for each LLM $i$, cognitive dimension $j$, and feedback condition $k$, we calculated the incremental effect as:
$\Delta Y_{ijk}=Y_{ijk}^{\text{(Feedback)}}-Y_{ij}^{\text{(Self-Awareness)}}.$
Feedback intensity was numerically encoded as Feedback-None $=0$, Feedback-Less $=1$, and Feedback-Most $=2$. Given the limited number of intensity levels, we restricted the analysis to linear and quadratic trends to prevent overfitting. The resulting polynomial trend model was defined as:
$\Delta Y_{ijk}=\mu+\alpha_{j}+\gamma_{1j}X_{k}+\gamma_{2j}X_{k}^{2}+s_{i}+\varepsilon_{ijk},$
where $X_{k}$ represents the numerical encoding of feedback intensity, and parameters $\gamma_{1j}$ and $\gamma_{2j}$ capture linear and quadratic trend effects for cognitive dimension $j$, respectively.
We performed separate hypothesis tests for each polynomial coefficient:
$H_{0}^{(\text{linear})}:\gamma_{1j}=0,\quad H_{1}^{(\text{linear})}:\gamma_{1j}\neq 0;$
$H_{0}^{(\text{quadratic})}:\gamma_{2j}=0,\quad H_{1}^{(\text{quadratic})}:\gamma_{2j}\neq 0.$
A significant linear trend ($\gamma_{1j}\neq 0$) indicates a monotonic incremental response to feedback intensity, whereas a significant quadratic trend ($\gamma_{2j}\neq 0$) captures more nuanced, nonlinear incremental dynamics, such as initially minimal changes followed by substantial shifts under higher-intensity feedback.
H.2 Results
Model Dim. Pair Compared Conditions $\mathbf{T}$ $\mathbf{df}$ $p_{\text{corr}}$
Qwen2.5-72B Fatal. SA–FL SA $>$ FL 6.90 19 $8.4\times 10^{-6}$
SA–FN SA $>$ FN 5.10 19 $3.8\times 10^{-4}$
DeepSeek-V3 Hier. FL–FM FL $<$ FM $-3.27$ 19 .024
Indiv. FN–SA FN $<$ SA $-4.70$ 19 .00093
Gemini-2.0 Egal. FM–FN FM $>$ FN 9.30 19 $1.0\times 10^{-7}$
Indiv. FM–FN FM $>$ FN 11.46 19 $3.4\times 10^{-9}$
Gemma-27B Indiv. FN–SA FN $<$ SA $-53.10$ 19 $2.4\times 10^{-21}$
GLM-9B Hier. SA–FM SA $<$ FM $-13.42$ 19 $2.3\times 10^{-10}$
GPT-4o Hier. FN–SA FN $<$ SA $-7.13$ 19 $4.4\times 10^{-6}$
InternLM-20B Fatal. FM–SA FM $>$ SA 39.91 19 $5.2\times 10^{-19}$
@@ H.2 Results
Model Dimension $\mathbf{F}$ $\mathbf{p}$ $\boldsymbol{\eta^{2}_{p}}$
Qwen2.5-72B Egalitarianism 3.35 .0250 .150 $\dagger$
Fatalism 23.02 $6.9\times 10^{-10}$548 $\dagger$
Hierarchy 1.26 .297 .062
Individualism 2.11 .109 .100
DeepSeek-V3 Egalitarianism 3.80 .0149 .167 $\dagger$
Fatalism 8.66 $7.9\times 10^{-5}$313 $\dagger$
Hierarchy 6.37 .00085 .251 $\dagger$
Individualism 21.99 $1.4\times 10^{-9}$536 $\dagger$
Gemini-2.0-Flash Egalitarianism 51.56 $3.0\times 10^{-16}$731 $\dagger$
Fatalism 20.27 $4.6\times 10^{-9}$516 $\dagger$
Hierarchy 43.69 $8.6\times 10^{-15}$697 $\dagger$
Individualism 105.77 $2.9\times 10^{-23}$848 $\dagger$
Gemma-3-27B-IT Egalitarianism 163.06 $6.2\times 10^{-28}$896 $\dagger$
Fatalism 8.09 .000141 .299 $\dagger$
Hierarchy 115.42 $3.4\times 10^{-24}$859 $\dagger$
Individualism 688.34 $1.0\times 10^{-44}$973 $\dagger$
GLM-4-9B Egalitarianism 59.93 $1.3\times 10^{-17}$759 $\dagger$
Fatalism 259.80 $3.3\times 10^{-33}$932 $\dagger$
Hierarchy 47.49 $1.6\times 10^{-15}$714 $\dagger$
Individualism 101.91 $7.0\times 10^{-23}$843 $\dagger$
GPT-4o Egalitarianism 15.86 $1.3\times 10^{-7}$455 $\dagger$
Fatalism 0.67 .574 .034
Hierarchy 30.46 $7.0\times 10^{-12}$616 $\dagger$
Individualism 53.95 $1.2\times 10^{-16}$740 $\dagger$
InternLM-20B Egalitarianism 33.70 $1.2\times 10^{-12}$639 $\dagger$
Fatalism 138.25 $4.0\times 10^{-26}$879 $\dagger$
Hierarchy 99.19 $1.3\times 10^{-22}$839 $\dagger$
Individualism 150.33 $4.9\times 10^{-27}$888 $\dagger$
Llama-3.3-70B Egalitarianism 26.78 $6.1\times 10^{-11}$585 $\dagger$
@@
Hierarchy 28.46 $2.2\times 10^{-11}$600 $\dagger$
Individualism 9.54 $3.4\times 10^{-5}$334 $\dagger$
Phi-4 Egalitarianism 16.80 $6.2\times 10^{-8}$469 $\dagger$
Fatalism 17.56 $3.4\times 10^{-8}$480 $\dagger$
@@
Individualism 12.87 $1.6\times 10^{-6}$404 $\dagger$
[…]
Figure 11: Statistically significant analysis results for every model-dimension trend. The Linear and Quadratic indicate two kinds of hypothetical trends. Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Model Dimension Linear ($r$) Quadratic ($r$)
Qwen2.5-72B-Instruct Egalitarianism 0.335 -0.194
Fatalism -0.202 -0.583∗∗∗
Hierarchy 0.161 -0.217
Individualism 0.274 -0.158
deepseek-chat Egalitarianism 0.430∗∗ 0.000
Fatalism 0.369 0.024
Hierarchy 0.491∗∗ 0.170
Individualism 0.755∗∗∗ 0.026
gemini-2.0-flash Egalitarianism 0.850∗∗∗ 0.080
Fatalism -0.458∗∗ -0.384
Hierarchy 0.833∗∗∗ 0.182
Individualism 0.896∗∗∗ 0.204
gemma-3-27b-it Egalitarianism 0.923∗∗∗ 0.188
Fatalism 0.528∗∗∗ -0.290
Hierarchy 0.898∗∗∗ 0.214
Individualism 0.953∗∗∗ 0.247
glm-4-9b-chat Egalitarianism 0.387 0.000
Fatalism 0.548∗∗∗ 0.000
Hierarchy 0.694∗∗∗ 0.189
Individualism 0.806∗∗∗ 0.114
gpt-4o Egalitarianism 0.635∗∗∗ -0.100
Fatalism 0.183 0.075
Hierarchy 0.748∗∗∗ -0.110
Individualism 0.830∗∗∗ -0.190
internlm2_5-20b-chat Egalitarianism 0.837∗∗∗ 0.183
Fatalism 0.895∗∗∗ 0.157
Hierarchy 0.916∗∗∗ 0.218
Individualism 0.903∗∗∗ 0.329
llama-3.3-70b-instruct Egalitarianism 0.769∗∗∗ 0.243
Fatalism 0.571∗∗∗ 0.033
Hierarchy 0.819∗∗∗ -0.049
Individualism 0.566∗∗∗ 0.245
phi-4 Egalitarianism 0.548∗∗∗ 0.211
Fatalism -0.441∗∗ -0.204
Hierarchy 0.056 0.226
Individualism 0.087 0.050
Table 17: Polynomial-trend results on incremental scores ($\Delta Y=Y_{\text{feedback}}-Y_{\text{self\mbox{-}aware}}$). Stars indicate significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$. Feedback intensity is coded as $\{-1,0,+1\}$ for None, Little, Most.
Model Dimension Linear ($r$) Quadratic ($r$)
Qwen2.5-72B-Instruct Egalitarianism 0.269 0.212
Fatalism $-0.356^{*}$ $-0.443^{**}$
Hierarchy 0.095 0.046
Individualism 0.219 0.173
deepseek-chat Egalitarianism $0.413^{**}$ $0.384^{*}$
Fatalism $0.362^{*}$ $0.340^{*}$
Hierarchy $0.519^{***}$ $0.515^{***}$
Individualism $0.733^{***}$ $0.686^{***}$
gemini-2.0-flash Egalitarianism $0.839^{***}$ $0.795^{***}$
Fatalism $-0.546^{***}$ $-0.582^{***}$
Hierarchy $0.851^{***}$ $0.826^{***}$
Individualism $0.917^{***}$ $0.891^{***}$
gemma-3-27b-it Egalitarianism $0.939^{***}$ $0.908^{***}$
Fatalism $0.427^{**}$ $0.341^{*}$
Hierarchy $0.922^{***}$ $0.898^{***}$
Individualism $0.985^{***}$ $0.962^{***}$
glm-4-9b-chat Egalitarianism $0.372^{*}$ $0.346^{*}$
Fatalism $0.526^{***}$ $0.489^{**}$
Hierarchy $0.719^{***}$ $0.704^{***}$
Individualism $0.806^{***}$ $0.770^{***}$
gpt-4o Egalitarianism $0.582^{***}$ $0.522^{***}$
Fatalism 0.197 0.197
Hierarchy $0.688^{***}$ $0.618^{***}$
Individualism $0.745^{***}$ $0.655^{***}$
internlm2_5-20b-chat Egalitarianism $0.855^{***}$ $0.830^{***}$
Fatalism $0.904^{***}$ $0.870^{***}$
Hierarchy $0.940^{***}$ $0.916^{***}$
Individualism $0.959^{***}$ $0.955^{***}$
llama-3.3-70b-instruct Egalitarianism $0.806^{***}$ $0.796^{***}$
Fatalism $0.558^{***}$ $0.525^{***}$
Hierarchy $0.773^{***}$ $0.709^{***}$
Individualism $0.611^{***}$ $0.615^{***}$
phi-4 Egalitarianism $0.585^{***}$ $0.584^{***}$
Fatalism $-0.481^{**}$ $-0.486^{**}$
Hierarchy 0.116 0.152
Individualism 0.097 0.100
Table 18: Polynomial–trend results on incremental scores ($\Delta Y=Y_{\text{feedback}}-Y_{\text{self\mbox{-}aware}}$). Significance: ${}^{*}\!p<0.05$, ${}^{**}\!p<0.01$, ${}^{***}\!p<0.001$. Feedback intensity is coded as $\{0,1,4\}$ for None, Little, Most.
Model Dimension Linear ($r$) Quadratic ($r$)
Qwen2.5-72B-Instruct Egalitarianism 0.335 0.269
Fatalism --0.202 --0.356
Hierarchy 0.161 0.095
Individualism 0.274 0.219
deepseek-chat Egalitarianism 0.430∗∗ 0.413∗∗
Fatalism 0.369 0.362
Hierarchy 0.491∗∗ 0.519∗∗∗
Individualism 0.755∗∗∗ 0.733∗∗∗
gemini-2.0-flash Egalitarianism 0.850∗∗∗ 0.839∗∗∗
Fatalism --0.458∗∗ --0.546∗∗∗
Hierarchy 0.833∗∗∗ 0.851∗∗∗
Individualism 0.896∗∗∗ 0.917∗∗∗
gemma-3-27b-it Egalitarianism 0.923∗∗∗ 0.939∗∗∗
Fatalism 0.528∗∗∗ 0.427∗∗
Hierarchy 0.898∗∗∗ 0.922∗∗∗
Individualism 0.953∗∗∗ 0.985∗∗∗
glm-4-9b-chat Egalitarianism 0.387 0.372
Fatalism 0.549∗∗∗ 0.526∗∗∗
Hierarchy 0.694∗∗∗ 0.719∗∗∗
Individualism 0.806∗∗∗ 0.806∗∗∗
gpt-4o Egalitarianism 0.635∗∗∗ 0.582∗∗∗
Fatalism 0.183 0.197
Hierarchy 0.748∗∗∗ 0.688∗∗∗
Individualism 0.830∗∗∗ 0.745∗∗∗
internlm2_5-20b-chat Egalitarianism 0.837∗∗∗ 0.855∗∗∗
Fatalism 0.895∗∗∗ 0.904∗∗∗
Hierarchy 0.916∗∗∗ 0.940∗∗∗
Individualism 0.903∗∗∗ 0.959∗∗∗
llama-3.3-70b-instruct Egalitarianism 0.769∗∗∗ 0.806∗∗∗
Fatalism 0.571∗∗∗ 0.558∗∗∗
Hierarchy 0.819∗∗∗ 0.773∗∗∗
Individualism 0.566∗∗∗ 0.611∗∗∗
phi-4 Egalitarianism 0.549∗∗∗ 0.585∗∗∗
Fatalism --0.441∗∗ --0.481∗∗
Hierarchy 0.056 0.116
Individualism 0.087 0.097
Table 19: Polynomial-trend correlations on incremental scores ($\Delta Y=Y_{\text{feedback}}-Y_{\text{self\mbox{-}aware}}$). Feedback intensity is coded as $\{0,1,2\}$ for None, Little, Most. Significance levels: ${}^{*}p<0.05$, ${}^{**}p<0.01$, ${}^{***}p<0.001$.
Appendix I Additional Results
@@
Figures 131315 and 15 illustrate the correlation of LLM Responses with Base Model by Family and Dimension under the baseline condition. Each bar represents the Pearson correlation between a smaller model and its base model within the same family, measured across one of four cognitive dimensions. The closer the value is to 1, the more similar the response profile. The plots reveal how reliably smaller-scale LLMs mimic the socio-cognitive outputs of their larger family counterparts. In the same way, Figures 171719 and 19 demonstrates the results under the self-awareness condition.
@@
Appendix J Prompts
In this appendix, we provide the exact prompts utilized in our experiments, including those for the automated questionnaire generation pipeline and for eliciting responses from large language models (LLMs). Prompts 1–5 (JJ) correspond to the distinct experimental conditions designed to test LLM responses under varying degrees of social referencing, as outlined in Section 4. Prompts 6–9 (JJ) detail the automated multi-agent prompting framework described in Appendix C, which generates, validates, and refines the questionnaire items for the Cognitive Attitude Dataset (CAD).
Prompt J.1: Questionnaire Answer Base Please answer the following question according to the provided instructions. Provide your rating along with a reason supporting your answer. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} Your response should strictly follow this JSON format:
@@
Prompt J.2: Questionnaire Answer Self-Awareness Please answer the following question according to the provided instructions. Include your rating and a reason for your decision. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Note that your response will be referenced by other humans and may influence their subsequent decisions. Question: {question} Your response should strictly follow this JSON format:
@@
Prompt J.3: Questionnaire Answer Feedback with Most Agreement Please answer the following question based on the provided instructions, considering the historical context provided below. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} History of the previous round: {history} In the previous round, other participants referenced your decision; 4 out of 5 participants agreed with your stance. Considering this information, please provide your updated answer. Your response should strictly follow this JSON format:
@@
Prompt J.4: Questionnaire Answer Feedback with Little Agreement Please answer the following question based on the provided instructions, considering the historical context provided below. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} History of the previous round: {history} In the previous round, other participants referenced your decision; only 1 out of 5 participants agreed with your stance. Considering this information, please provide your updated answer. Your response should strictly follow this JSON format:
@@
Prompt J.5: Questionnaire Answer Feedback with No Agreement Please answer the following question based on the provided instructions, considering the historical context provided below. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} History of the previous round: {history} In the previous round, other participants referenced your decision; none of the 5 participants agreed with your stance. Considering this information, please provide your updated answer. Your response should strictly follow this JSON format:
@@
Prompt J.6: Generate Questionnaire You are now generating a structured questionnaire to measure worldview according to the provided taxonomy and specifically targeting the designated sub-dimension. Maintain a clear mental map of all dimensions and sub-dimensions in the provided taxonomy. For the given sub-dimension, generate exactly 20 distinct Likert-scale questions. Each question must clearly relate to the targeted sub-dimension and must not duplicate or overlap with any other question. Each question should be designed for evaluation using a Likert scale (1 = Strongly Disagree, 5 = Strongly Agree). Sub-dimension: {target_subdimension} Taxonomy: {taxonomy} Output Format: Output your response strictly in the following JSON format:
@@
Prompt J.7: Validate Adherence You are now tasked with evaluating the adherence of each question in the provided questionnaire to the specified sub-dimension within the given taxonomy. Maintain a clear mental map of all dimensions and sub-dimensions in the provided taxonomy. Review each question carefully, assessing how strictly it aligns with the targeted sub-dimension. For each question, assign an adherence score from 1 (very weak adherence) to 5 (very strong adherence) along with a brief justification for your rating. Provided Information: Questionnaire: {questionnaire (JSON format)} Taxonomy: {taxonomy (JSON format)} Sub-dimension: {target_subdimension} Output Format: Output your response strictly in the following JSON format:
@@
Prompt J.8: Validate Measurability You are now tasked with evaluating the measurable categorization of each question within the provided questionnaire. Each question is intended to be assessed using a Likert scale (1 = Strongly Disagree, 5 = Strongly Agree). Carefully review each question and judge whether it can be correctly and effectively measured using the specified Likert scale. For each question, assign: Measure: 1 if the question can be clearly and effectively measured using the Likert scale provided, or 0 if it cannot. Reason: Provide a brief explanation supporting your evaluation. Provided Information: Questionnaire: {questionnaire (JSON format)} Output Format: Output your response strictly in the following JSON format:
@@
Prompt J.9: Refine Questionnaire You are now tasked with refining questions from the provided questionnaire based on evaluations of their (1) adherence to a specific sub-dimension within the given taxonomy, and (2) measurability using the Likert scale (1 = Strongly Disagree, 5 = Strongly Agree). Specifically, you must refine any question if: The adherence score to the targeted sub-dimension is lower than 3, or The measure score is 0 (indicating the question is not measurable effectively). Refine each identified question to ensure it clearly aligns with the targeted sub-dimension and can be effectively measured using the specified Likert scale. Only include questions that require refinement. Provided Information: Questionnaire: {questionnaire (JSON format)} Taxonomy: {taxonomy (JSON format)} Sub-dimension: {target_subdimension} Adherence Evaluation: {adherence_evals (JSON format)} Measurability Evaluation: {measurability_evals (JSON format)} Output Format: Output your response strictly in the following JSON format:

### v2

@@ F.1 Results
DeepSeek‑V3 Egalitarianism 3.80 .0149 .167 $\dagger$
Fatalism 8.66 $7.9\times 10^{-5}$ .313 $\dagger$
Hierarchy 6.37 .00085 .251 $\dagger$
Individualism 21.99 $1.4\times 10^{-9}$ .536 $\dagger$
Gemini‑2.0‑Flash Egalitarianism 51.56 $3.0\times 10^{-16}$ .731 $\dagger$
Fatalism 20.27 $4.6\times 10^{-9}$ .516 $\dagger$
Hierarchy 43.69 $8.6\times 10^{-15}$ .697 $\dagger$
Individualism 105.77 $2.9\times 10^{-23}$ .848 $\dagger$
Gemma‑3‑27B‑IT Egalitarianism 163.06 $6.2\times 10^{-28}$ .896 $\dagger$
Fatalism 8.09 .000141 .299 $\dagger$
Hierarchy 115.42 $3.4\times 10^{-24}$ .859 $\dagger$
Individualism 688.34 $1.0\times 10^{-44}$ .973 $\dagger$
GLM‑4‑9B Egalitarianism 59.93 $1.3\times 10^{-17}$ .759 $\dagger$
Fatalism 259.80 $3.3\times 10^{-33}$ .932 $\dagger$
Hierarchy 47.49 $1.6\times 10^{-15}$ .714 $\dagger$
Individualism 101.91 $7.0\times 10^{-23}$ .843 $\dagger$
GPT‑4o Egalitarianism 15.86 $1.3\times 10^{-7}$ .455 $\dagger$
@@
Hierarchy 30.46 $7.0\times 10^{-12}$ .616 $\dagger$
Individualism 53.95 $1.2\times 10^{-16}$ .740 $\dagger$
InternLM‑20B Egalitarianism 33.70 $1.2\times 10^{-12}$ .639 $\dagger$
Fatalism 138.25 $4.0\times 10^{-26}$ .879 $\dagger$
Hierarchy 99.19 $1.3\times 10^{-22}$ .839 $\dagger$
Individualism 150.33 $4.9\times 10^{-27}$ .888 $\dagger$
Llama‑3.3‑70B Egalitarianism 26.78 $6.1\times 10^{-11}$ .585 $\dagger$
@@
Hierarchy 28.46 $2.2\times 10^{-11}$ .600 $\dagger$
Individualism 9.54 $4\times 10^{-5} .33$4 $\dagger$
Phi‑4 Egalitarianism 16.80 $6.2\times 10^{-8}$ .469 $\dagger$
Fatalism 17.56 $3.4\times 10^{-8}$ .480 $\dagger$
@@
Individualism 12.87 $1.6\times 10^{-6}$ .404 $\dagger$
Appendix G Additional Results
@@
Figures 557 and 7 illustrate the correlation of LLM Responses with Base Model by Family and Dimension under the baseline condition. Each bar represents the Pearson correlation between a smaller model and its base model within the same family, measured across one of four cognitive dimensions. The closer the value is to 1, the more similar the response profile. The plots reveal how reliably smaller-scale LLMs mimic the socio-cognitive outputs of their larger family counterparts. In the same way, Figures 9911 and 11 demonstrates the results under the self-awareness condition.
@@
Appendix H Prompts
In this appendix, we provide the exact prompts utilized in our experiments, including those for the automated questionnaire generation pipeline and for eliciting responses from large language models (LLMs). Prompts 1–5 ($H$) correspond to the distinct experimental conditions designed to test LLM responses under varying degrees of social referencing. Prompts 6–9 ($H$) detail the automated multi-agent prompting framework described in Appendix A, which generates, validates, and refines the questionnaire items for the Cognitive Attitude Dataset (CAD).
Prompt H.1: Questionnaire Answer Base Please answer the following question according to the provided instructions. Provide your rating along with a reason supporting your answer. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} Your response should strictly follow this JSON format:
@@ Appendix H Prompts
Prompt H.2: Questionnaire Answer Self-Awareness Please answer the following question according to the provided instructions. Include your rating and a reason for your decision. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Note that your response will be referenced by other humans and may influence their subsequent decisions. Question: {question} Your response should strictly follow this JSON format:
@@
Prompt H.3: Questionnaire Answer Feedback with Most Agreement Please answer the following question based on the provided instructions, considering the historical context provided below. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} History of the previous round: {history} In the previous round, other participants referenced your decision; 4 out of 5 participants agreed with your stance. Considering this information, please provide your updated answer. Your response should strictly follow this JSON format:
@@
Prompt H.4: Questionnaire Answer Feedback with Little Agreement Please answer the following question based on the provided instructions, considering the historical context provided below. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} History of the previous round: {history} In the previous round, other participants referenced your decision; only 1 out of 5 participants agreed with your stance. Considering this information, please provide your updated answer. Your response should strictly follow this JSON format:
@@
Prompt H.5: Questionnaire Answer Feedback with No Agreement Please answer the following question based on the provided instructions, considering the historical context provided below. Instruction: Please indicate your agreement or disagreement using a Likert scale (1=Strongly Disagree, 5=Strongly Agree). Question: {question} History of the previous round: {history} In the previous round, other participants referenced your decision; none of the 5 participants agreed with your stance. Considering this information, please provide your updated answer. Your response should strictly follow this JSON format:
@@
Prompt H.6: Generate Questionnaire You are now generating a structured questionnaire to measure worldview according to the provided taxonomy and specifically targeting the designated sub-dimension. Maintain a clear mental map of all dimensions and sub-dimensions in the provided taxonomy. For the given sub-dimension, generate exactly 20 distinct Likert-scale questions. Each question must clearly relate to the targeted sub-dimension and must not duplicate or overlap with any other question. Each question should be designed for evaluation using a Likert scale (1 = Strongly Disagree, 5 = Strongly Agree). Sub-dimension: {target_subdimension} Taxonomy: {taxonomy} Output Format: Output your response strictly in the following JSON format:
@@
Prompt H.7: Validate Adherence You are now tasked with evaluating the adherence of each question in the provided questionnaire to the specified sub-dimension within the given taxonomy. Maintain a clear mental map of all dimensions and sub-dimensions in the provided taxonomy. Review each question carefully, assessing how strictly it aligns with the targeted sub-dimension. For each question, assign an adherence score from 1 (very weak adherence) to 5 (very strong adherence) along with a brief justification for your rating. Provided Information: Questionnaire: {questionnaire (JSON format)} Taxonomy: {taxonomy (JSON format)} Sub-dimension: {target_subdimension} Output Format: Output your response strictly in the following JSON format:
@@
Prompt H.8: Validate Measurability You are now tasked with evaluating the measurable categorization of each question within the provided questionnaire. Each question is intended to be assessed using a Likert scale (1 = Strongly Disagree, 5 = Strongly Agree). Carefully review each question and judge whether it can be correctly and effectively measured using the specified Likert scale. For each question, assign: Measure: 1 if the question can be clearly and effectively measured using the Likert scale provided, or 0 if it cannot. Reason: Provide a brief explanation supporting your evaluation. Provided Information: Questionnaire: {questionnaire (JSON format)} Output Format: Output your response strictly in the following JSON format:
@@
Prompt H.9: Refine Questionnaire You are now tasked with refining questions from the provided questionnaire based on evaluations of their (1) adherence to a specific sub-dimension within the given taxonomy, and (2) measurability using the Likert scale (1 = Strongly Disagree, 5 = Strongly Agree). Specifically, you must refine any question if: The adherence score to the targeted sub-dimension is lower than 3, or The measure score is 0 (indicating the question is not measurable effectively). Refine each identified question to ensure it clearly aligns with the targeted sub-dimension and can be effectively measured using the specified Likert scale. Only include questions that require refinement. Provided Information: Questionnaire: {questionnaire (JSON format)} Taxonomy: {taxonomy (JSON format)} Sub-dimension: {target_subdimension} Adherence Evaluation: {adherence_evals (JSON format)} Measurability Evaluation: {measurability_evals (JSON format)} Output Format: Output your response strictly in the following JSON format: